OpenStudy (anonymous):

A membership to Movie Night Movie Club costs $10, plus $2 per movie. Make a T-table for the equation for this problem, using the x-values 0, 5, 10. Type in the ordered pairs you get as your answer.

OpenStudy (anonymous):

do u want help or just answers

hero (hero):

10 + 2(0) = ? 10 + 2(5) = ? 10 = 2(10) = ?

hero (hero):

I think he might want both

OpenStudy (anonymous):

but hero buddy here ppl are interested in answers only

hero (hero):

Apparently

OpenStudy (anonymous):

see here u have given the hint but still no answer but no response........

hero (hero):

I did get a response. He medaled me

OpenStudy (anonymous):

hahahaha i too had medaled u does it mean i dint knew the answer!!!!!!!!!

hero (hero):

If he had any further questions, he would ask.

OpenStudy (anonymous):

hmmmm but u will be happy if it will be posted <------------somewhere here right.......